OpenStudy (anonymous):

Hey are you up for a challenge? Please show all work:) 2x-6/4 = x-3/5 A jacket at Khols cost $80. The jacket was put on sale for $64. What was the percent of change of the cost in this jacket. What percent of 25 is 300? This one is a long one:) A survey of 200 people said that 30% of those surveyed enjoyed reading a book over watching television. If the survey taker asks 50 more people, and 40 of them say they enjoy reading over tv, what is the NEW percent of people who enjoy reading over tv?

OpenStudy (anonymous):

Easy

OpenStudy (anonymous):

Yeah the "challenge" aspect was the fact that there were multiple questions.

OpenStudy (anonymous):

Problem #1 2x-6/4 = x-3/5 subtract x from both sides x - 6/4 = -3/5 add 1.5 to both sides x = - 3/5 + 1.5 -3/5 is -.6 so x = .9

OpenStudy (anonymous):

Problem #2 Jacket = $80 SalePrice = $64 Discount is defined as Jacket - SalePrice = $16 % of the discount is Discount / Jacket = 20% answer = %20

OpenStudy (anonymous):

Problem #3 To find what percent 300 is of 25, divide 300 into groups of 25. 4 X 25 = 100 but we want 3 times that amount so multiply both sides by 3 12 X 25 = 300 to express the growth as a percentage, take %100 of 25 = 25 and multiply both sides by 12 12 X %100 X 25 = 300 answer = %1200

OpenStudy (anonymous):

Problem #4 Original Study Group = 200 people %30 are readers therefore 200 people X .3 = 60 people (don't forget to explicitly write the units down) Extended Group = 50 more people The reader subset is 40 out of 50 or .8 To extend the study we need to add 40 new readers to the 60 readers in the original study. Then divide by the total number of people surveyed. (40+60) / 250 = .4 or %40